# EmoBipedNav: Emotion-aware Social Navigation for Bipedal Robots with Deep Reinforcement Learning

### v1

@@ Abstract
This study presents an emotion-aware navigation framework – EmoBipedNav – using deep reinforcement learning (DRL) for bipedal robots walking in socially interactive environments. The inherent locomotion constraints of bipedal robots challenge their safe maneuvering capabilities in dynamic environments. When combined with the intricacies of social environments, including pedestrian interactions and social cues, such as emotions, these challenges become even more pronounced. To address these coupled problems, we propose a two-stage pipeline that considers both bipedal locomotion constraints and complex social environments. Specifically, social navigation scenarios are represented using sequential LiDAR grid maps (LGMs), from which we extract latent features, including collision regions, emotion-related discomfort zones, social interactions, and the spatio-temporal dynamics of evolving environments. The extracted features are directly mapped to the actions of reduced-order models (ROMs) through a DRL architecture. Furthermore, the proposed framework incorporates full-order dynamics and locomotion constraints during training, effectively accounting for tracking errors and restrictions of the locomotion controller while planning the trajectory with ROMs. Comprehensive experiments demonstrate that our approach exceeds both model-based planners and DRL-based baselines. The hardware videos and open-source code are available at https://gatech-lidar.github.io/emobipednav.github.io/.
Index Terms:
Bipedal robot, locomotion, social navigation, emotion-aware navigation, deep reinforcement learning.
@@ I INTRODUCTION
  • Integration of pedestrian emotions. Our framework incorporates pedestrian emotions to design discomfort zones and further enable emotion-aware navigation.
@@
  • Comprehensive benchmarks and hardware evaluations. We either reproduce or improve SOTA benchmarks for bipedal social navigation, validating our pipeline against both model-based planners and DRL-based baselines. In addition, we successfully transfer the simulated navigation policy to the real bipedal robot Digit.
@@ III-A Preliminaries
Given the control $\mathbf{u}_{k}$, the CoM displacement $\Delta x^{\text{loc}}$, and the LIP model, the CoM position is propagated from $p_{k}$ to $p_{k+1}$, and the sagittal speed changes to $\upsilon_{k+1}^{\text{loc}}$ from $\upsilon_{k}^{\text{loc}}$. However, there are discrepancies between the actual CoM position $p_{k+1}^{*}$ and $p_{k+1}$, and between the actual sagittal speed $*\upsilon_{k+1}^{\text{loc*}}$ and $\upsilon_{k+1}^{\text{loc}}$. More specifically, the full-body controller generates joint torques to track the desired CoM displacement $\Delta x^{\text{loc}}$ and the control $\mathbf{u}_{k}$, as illustrated in Fig. 2-(d). However, discrepancies between the LIP model and the full-body dynamics can cause the robot’s states to deviate from the desired ones, which may lead to collisions and intrusions into discomfort areas of pedestrians, as demonstrated in Section IV. Furthermore, due to the limitations of Digit’s hardware and locomotion controllers, these motion constraints are significantly more stringent than those of pedestrians, making Digit less reactive to pedestrians with high walking speed (e.g., $1$ m/s), thus posing additional challenges for Digit navigating through crowded pedestrian environments. In Section IV, we show the benchmark results to demonstrate that these constraints significantly degrade the navigation performance of the SOTA approaches.
@@ III-C Observation and State Spaces
Alternatively, occupancy grid maps (OGMs)  [^1] http://wiki.ros.org/costmap-2d, as shown in Fig. 3-(c), are commonly used for local collision avoidance [41]. With the same map size $M^{2}$ and detection radius $L$, the resolution of each grid is calculated evenly as $(M/2L)^{2}$. In contrast, the resolution of our LGM is calculated on average as $(M/\sqrt{\pi}L)^{2}$, demonstrating that our LGM representation has a higher resolution on average. More specifically, the LGM has a higher resolution for grids closer to the map origin and a lower resolution for grids farther away. For example, when $M=100$, $L=6.0$ m, and $K=1800$, for $i=1$, the LGM has 100 grids while the OGM only has 4. When $i>13$, the density of the OGM grid surpasses that of the LGM. As a result, the LGM can represent more details within the $0.7813\cdot L/M=$ 0.78 m range of the ego-agent, as illustrated in the top figures of Fig. 3. Given the same obstacles close to the ego-agent, their representation in Fig. 3-(b) (LGM) contains more pixels compared to Fig. 3-(c) (OGM), indicating that LGM can capture more details when obstacles are near, which is beneficial for collision avoidance and alleviating discomfort, as demonstrated in Section IV.
@@
The observation $\mathbf{o}_{k}$ at the current walking step is the LGM capturing only the static immediate environment in the spatial dimension, which might encourage conservative motions to avoid collisions as quantitatively analyzed in Section IV. To capture dynamic information in the temporal dimension, such as pedestrian-pedestrian and pedestrian-robot interactions, and the evolving dynamics of the navigation system, we incorporate previous LGMs from time step $k-N+1$ to $k-1$ as illustrated in Fig. 2-(a). Thus, the observation in both the spatial and temporal dimensions is represented as $\mathbf{o}_{k}^{\rm{sp-tem}}=\{\mathbf{o}_{k-N+1},...,\mathbf{o}_{k-1},\mathbf{o}_{k}\}$. From spatio-temporal LGMs, we can extract latent features, including collision areas, complex pedestrian-pedestrian and pedestrian-robot interactions, emotion-related discomfort zones, and the dynamic evolution of the navigation system.
@@ III-E Reward Integration
$r_{k}^{\rm{col}}=\begin{cases}-0.6&\text{if}~{}d^{\rm{all}}_{\min}\leq r^{\rm{robot}},\\ -0.1&\text{else if}~{}d^{\rm{static}}_{\min}<d^{\rm{static}}+r^{\rm{robot}},\\ 0.0&\text{else},\\ \end{cases}$ (7)
@@
$r_{k}^{\rm{goal}}=\begin{cases}0.5&\text{if}~{}d_{k}^{\rm{goal}}\leq d^{\rm{goal}},\\ 0.3\cdot(d_{k-1}^{\rm{goal}}-d_{k}^{\rm{goal}})&\text{else},\\ \end{cases}$ (8)
@@
$r_{k}^{\rm{emo}}=\begin{cases}-0.1&\text{if}~{}\mathcal{C},\\ 0.0&\text{else},\\ \end{cases}$ (9)
@@ IV Experiments
$N$ 9 $d^{\rm{goal}}$ $0.1$ m $\bar{u}^{\Delta\theta}$ $0.2$ rad
@@
We first detail the parameter specifications and the network architecture, followed by the environmental settings. Next, we benchmark our proposed navigation approach against SOTA methods that do not account for emotion-related observations and rewards. In addition, we implement another training with emotion-related observations and reward functions. Further investigations on time-varying emotions are carried out. Finally, we demonstrate the deployment of our navigation policy on the bipedal robot Digit.
@@ IV-B Environment Settings
Static obstacles. Static obstacles are randomly placed within a rectangular area, with the $x$-coordinate ranging from $-3$ to $3$ m and the $y$-coordinate from $-1.5$ to $1.5$ m. Each obstacle has a radius ranging from $0.2$ to $0.4$ m.
@@
Robot. The robot navigates from position $(-4,0)$ to position $(4,0)$ m, interacting with pedestrians and actively avoiding static obstacles. We simulate the bipedal robot Digit in MuJoCo [42]. The navigation policy in the high-level motion planner only generates the sagittal speed and heading angle derived from the ROM.
@@
IV-C Benchmarks
@@
IV-D Emotion-integrated Implementation
@@
Var-DD$=[0.2,0.35,0.5]$ m $\leftrightarrow$ [happy, neutral, negative]
@@
IV-E Implementation with Time-varying Emotions
@@ IV-F Hardware Demonstrations
[…]

### v2

@@ Abstract
This study presents an emotion-aware navigation framework – EmoBipedNav – using deep reinforcement learning (DRL) for bipedal robots walking in socially interactive environments. The inherent complex dynamics of bipedal robots challenge their safe maneuvering capabilities in dynamic environments. Furthermore, the intricacies of social interactions and cues such as emotions significantly compound the challenges of bipedal robot navigation. To address these coupled issues, we propose a two-stage pipeline that jointly considers the bipedal full-body dynamics and the complexities of socially aware navigation. More specifically, an emotion-integrated navigation policy is developed to balance safety, efficiency, and social courtesy by responding to human emotional cues. One key component of the policy is a novel representation of social environments using sequential LiDAR grid maps (LGMs), from which we extract latent features, implicitly including collision regions, discomfort zones determined by emotions, social interactions, and the evolving dynamics of the navigation system including robot movements and pedestrian motions. To accounts for path tracking errors and locomotion constraints during social navigation, we present an end-to-end navigation system that incorporates full-order robot dynamics during training. Finally, extensive benchmarking and sim-to-real experiments demonstrate that our method outperforms both model-based planners and DRL baselines, and generalizes effectively to real-world environments. The hardware videos and open-source code are available at https://gatech-lidar.github.io/emobipednav.github.io/.
@@ I INTRODUCTION
  • Emotion integration in socially interactive environments. Our framework incorporates pedestrian emotions and constructs various discomfort zones, enabling emotion-aware navigation in socially interactive environments.
@@
  • Benchmarks and hardware evaluations. We compare the performance of our method with SOTA baselines. In addition, we transfer the simulated navigation policy to the real bipedal robot Digit.
@@ III-A Preliminaries
Given the control $\mathbf{u}_{k}$, the CoM displacement $\Delta x^{\text{loc}}$, and the LIP model, the CoM position is propagated from $p_{k}$ to $p_{k+1}$, and the sagittal speed changes to $\upsilon_{k+1}^{\text{loc}}$ from $\upsilon_{k}^{\text{loc}}$. However, there are discrepancies between the actual CoM position $p_{k+1}^{*}$ and $p_{k+1}$, and between the actual sagittal speed $\upsilon_{k+1}^{\text{loc*}}$ and $\upsilon_{k+1}^{\text{loc}}$. More specifically, the full-body controller generates joint torques to track the desired CoM displacement $\Delta x^{\text{loc}}$ and the control $\mathbf{u}_{k}$, as illustrated in Fig. 2-(d). However, discrepancies between the LIP model and the full-body dynamics can cause the robot’s states to deviate from the desired ones, which may lead to collisions and intrusions into discomfort areas of pedestrians, as demonstrated in Section IV. Furthermore, due to the limitations of Digit’s hardware and locomotion controllers, these motion constraints are significantly more stringent than those of pedestrians, making Digit less reactive to pedestrians with high walking speed (e.g., $1$ m/s), thus posing additional challenges for Digit navigating through crowded pedestrian environments. In Section IV, we show the benchmark results to demonstrate that these constraints significantly degrade the navigation performance of the SOTA approaches.
@@ III-C Observation and State Spaces
Remark 2. During training, we assume pedestrian emotions are pre-known and randomly assign them to facilitate sim-to-real transfer. These emotions are used to define corresponding discomfort zones, which contribute to the construction of the LiDAR grid maps. In real-world implementations, pedestrian emotions are estimated from facial cues using a pre-trained CNN model  [^1] https://github.com/susantabiswas/realtime-facial-emotion-analyzer consisting of six convolutional layers and four fully connected layers.
@@
Alternatively, occupancy grid maps (OGMs)  [^2] http://wiki.ros.org/costmap-2d, as shown in Fig. 3-(c), are commonly used for local collision avoidance [41]. With the same map size $M^{2}$ and detection radius $L$, the resolution of each grid is calculated evenly as $(M/2L)^{2}$. In contrast, the resolution of our LGM is calculated on average as $(M/\sqrt{\pi}L)^{2}$, demonstrating that our LGM representation has a higher resolution on average. More specifically, the LGM has a higher resolution for grids closer to the map origin and a lower resolution for grids farther away. For example, when $M=100$, $L=6.0$ m, and $K=1800$, for $i=1$, the LGM has 100 grids while the OGM only has 4. When $i>13$, the density of the OGM grid surpasses that of the LGM. As a result, the LGM can represent more details within the $13\cdot L/M=0.78$ m range of the ego-agent, as illustrated in the top figures of Fig. 3. Given the same obstacles close to the ego-agent, their representation in Fig. 3-(b) (LGM) contains more pixels compared to Fig. 3-(c) (OGM), indicating that LGM can capture more details when obstacles are near, which is beneficial for collision avoidance and alleviating discomfort, as demonstrated in Section IV.
@@
The observation $\mathbf{o}_{k}$ at the current walking step is the LGM capturing only the static immediate environment in the spatial dimension, which might encourage conservative motions to avoid collisions as quantitatively analyzed in Section IV. To capture dynamic information in the temporal dimension, such as pedestrian-pedestrian and pedestrian-robot interactions, and the evolving dynamics of the navigation system, we incorporate previous LGMs from time step kN+1k-N+1 to k1k-1 as illustrated in Fig. 2-(a). Thus, the observation in both the spatial and temporal dimensions is represented as $\mathbf{o}_{k}^{\rm{sp-tem}}=\{\mathbf{o}_{k-N+1},...,\mathbf{o}_{k-1},\mathbf{o}_{k}\}$. From spatio-temporal LGMs, we can extract latent features, including collision areas, complex pedestrian-pedestrian and pedestrian-robot interactions, emotion-related discomfort zones, and the dynamic evolution of the navigation system.
@@ III-E Reward Integration
[…]
@@
[…]
@@
where $\mathcal{C}$ is a distance condition with respect to pedestrian emotions. Specifically, the discomfort distance of the $i^{\rm th}$ pedestrian is defined as $d_{p_{i}}^{\rm{emo}}\in\{0.2,0.35,0.5\}$m, depending on the corresponding emotion of the pedestrian $\{\texttt{happy},~\texttt{neutral},~\texttt{negative}\}$. $r^{\rm{ped}}$ is a constant pedestrian radius. Given the distance $d_{p_{i}}$ from the margin of the $i^{\rm th}$ pedestrian to the ego-agent, the penalty is 0.1-0.1 if $\exists~i$, such that $d_{p_{i}}<d_{p_{i}}^{\rm{emo}}+r^{\rm{robot}}$. Note that, to avoid learning a conservative navigation policy, we only count once if the ego-agent enters the discomfort zones of several objects, including static obstacles and dynamic pedestrians, at the same walking step. That means we add 0.1-0.1 only once in Eq. (6) although it appears in both Eq. (7) and Eq. (9).
Remark 3. Identifying the optimal reward configuration is challenging due to the highly coupled relationships among different reward terms. For example, when larger penalty factors were applied for entering discomfort zones (e.g., -0.6), the number of discomfort events decreased; however, this came at the cost of increased navigation time, as the agent adopted more conservative behaviors. Due to space limitations, we have chosen not to include the detailed results of these parameter variations. Instead, we empirically fine-tuned the reward function and ultimately selected a feasible and sub-optimal combination that balances safety and efficiency.
Our algorithm is summarized in Algorithm 1.
Algorithm 1 EmoBipedNav Algorithm
1:Initialize: SAC replay buffer $\mathcal{D}$, $k\leftarrow 1$
2:Reset: pedestrian positions and emotions, positions and sizes of static obstacles, robot position, and $\mathbf{u}_{k-1}\leftarrow\mathbf{0}$
3:Observe: $\mathbf{o}_{k}^{\rm{sp-tem}}$, $\mathbf{s}_{k}^{\rm{goal}}$, and $\mathbf{u}_{k-1}$
4:while $k\leq\mathcal{K}$ do
5:  Sample action $\mathbf{u}_{k}$ using Eq. (5a) and (5c)
6:  Compute the desired foot placement with ALIP [39] given $u_{k}^{\upsilon}\in\mathbf{u}_{k}$
7:  Track the full-body trajectory planned based on the foot placement using a passivity-based controller [44]
8:  Rotate hip yaw-joint to track heading angle $u_{k}^{\Delta\theta}\in\mathbf{u}_{k}$
9:  Joint torque control for Digit
10:  $k\leftarrow k+1$
11:  Observe
12:  Calculate reward with Eq. (6)-(9)
13:  if goal reached or collision or timeout then
14:   Reset
15:  end if
16:  Add observation, reward, and reset information to $\mathcal{D}$
17:  Update networks Eq. (5) using samples from $\mathcal{D}$ and SAC Algorithm
18:end while
@@ IV Experiments
NN 9 $d^{\rm{goal}}$ $0.1$ m $\bar{u}^{\Delta\theta}$ $0.2$ rad
@@
We first detail the parameter specifications and the network architecture, followed by the environmental settings. Subsequently, we experimentally evaluate the tracking errors introduced by the use of ROMs. Next, we benchmark our proposed navigation approach against SOTA methods that do not account for emotion-related observations and rewards. In addition, we implement another training with emotion-related observations and reward functions. Further investigations on time-varying emotions are carried out. Finally, we demonstrate the deployment of our navigation policy on the bipedal robot Digit.
@@ IV-B Environment Settings
Static obstacles. Static obstacles are randomly placed within a rectangular area, with the $x$-coordinate ranging from 3-3 to $3$ m and the $y$-coordinate from 1.5-1.5 to $1.5$ m. Each obstacle has a radius ranging from $0.2$ to $0.4$ m.
@@
Robot. The robot navigates from position (4,0)(-4,0) to position $(4,0)$ m, interacting with pedestrians and actively avoiding static obstacles. We simulate the bipedal robot Digit in MuJoCo [42]. The navigation policy in the high-level motion planner only generates the sagittal speed and heading angle derived from the ROM.
@@
IV-C Model Discrepancies
We describe the model discrepancies between the ROM and full-body bipedal dynamics in Section III-A. To clearly and intuitively illustrate these model gaps, we conduct simulation experiments that evaluate the tracking performance.
Figure 5: Tracking performance. The top figure shows the $x$ tracking, the middle plots the $y$ tracking, and the bottom illustrates the heading angle following in the world frame.
We select a representative simulation trial in MuJoCo, where Digit navigates a pedestrian-rich environment. For each walking step, we plot the desired $x$ and $y$ positions and the heading angle, alongside the actual positions and heading angle of the simulated Digit. The tracking results are shown in Fig. 5. While the tracking of the $x$-position is highly accurate, the $y$-position tracking error is non-negligible, and the heading angle exhibits more pronounced deviations. As discussed in Eq. (1), the ROM is primarily designed for sagittal motion and does not explicitly account for body orientation and lateral motion. As a result, $x$-direction tracking remains precise, whereas tracking errors in $y$-position and heading angle become more noticeable, particularly when the desired heading angle changes frequently. Such tracking errors significantly degrade navigation performance when deploying policies trained with ROMs onto full-body bipedal robots, as will be experimentally demonstrated in Section IV-E.
IV-D Benchmarks
@@
IV-E Emotion-integrated Implementation
@@
Var-DD=[0.2,0.35,0.5]=[0.2,0.35,0.5] m $\leftrightarrow$ [happy, neutral, negative]
@@
IV-F Implementation with Time-varying Emotions
@@ IV-G Hardware Demonstrations
Deployments integrating emotion recognition. To further validate the practicality of our pipeline with integrated pedestrian emotion recognition, we equip the physical Digit with a ZED stereo camera. Robot localization and pedestrian detection are performed using camera APIs  [^4] https://www.stereolabs.com/docs. For each detected pedestrian, facial features are extracted using OpenCV libraries  [^5] https://docs.opencv.org/3.4/db/d28/tutorial-cascade-classifier.html, and emotions are recognized using a pre-trained model 3. If no face is detected – due to factors such as long distance, occlusion by other pedestrians, or the pedestrian facing away from the camera – the emotion is set to neutral. Localization and pedestrian detection are performed on a Jetson Orin at 10 Hz, while face extraction and emotion recognition operate at 0.5 Hz. Meanwhile, the low-level controller and social navigation policy run on an Intel NUC at 1000 Hz and 2.5 Hz, respectively.